OpenStudy (anonymous):

Derivative Problem: If you're finding the derivative of t^3 isn't it 3t^3 instead of 3t^2 because you take t^3 = t(3t^2) so when you foil it 3t^2 * t = 3t^3 but if I'm wrong can someone please explain (:

OpenStudy (anonymous):

if you did that you would have to use the product rule so it is 3t^2

OpenStudy (anonymous):

are you familiar with the product rule?

OpenStudy (anonymous):

is it 3t^2 because when you write it out it's like 1(3t^2) instead of t(3t^2)?

OpenStudy (anonymous):

And I just learned the product rule last class, so I'm still eehhh about it.

OpenStudy (anonymous):

well, personally I would just never pull out the t. I would leave it as t^3 , there would be no need to pull out a t. so (d/dt)t^3=3t^2

OpenStudy (anonymous):

if you wanted to do the product rule of t*t^2 it would be (d/dt t)(t^2)+(t)(d/dt t^2)

OpenStudy (anonymous):

so 1(t^2)+t(2t)

OpenStudy (anonymous):

or t^2 +2t^2 = 3t^2 .. notice how we get the same answer.

OpenStudy (turingtest):

are you trying to apply the product rule or the chain rule here?

OpenStudy (anonymous):

that was just the product rule. you don't need either.

OpenStudy (turingtest):

If it's the product rule matt already covered it if it's the chain rule, then note that the inner function is t and d/dt(t)=1 so it would be d/dt(t^3)=3t^2(1)=3t^2 so as matt said, neither are necessary

OpenStudy (anonymous):

it was the product rule.